40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17138     Entry time: Tue Sep 13 14:12:03 2022     Reply to this: 17142
Author: Yehonathan 
Type: Update 
Category: BHD 
Subject: Trying LO phase locking again 

[Paco, Yehonathan]

Summary:  We locked LO phase using the DC PD (A - B) error point without saturating the control point, i.e. not a "bang bang" control.


Some suspensions were improved so we figure we should go back to trying to lock the LO phase.

We misalign ETMs and lock MICH using AS55. We put a small MICH offset by putting C1:LSC-MICH_OFFSET = -80.

AS and LO beams were aligned to overlap by maximizing the BHD signal visibility.

BHD DCPDs were balanced by misaligning the AS beam and using the LO beam only.

We measure the transfer function between the DCPDs and find the coherence is 1 at 1 Hz (because of seismic motion) so we measure the ratio between them to be 0.3db.

AS beam is aligned again to overlap with the LO beam. For the work below, we use the largest MICH OFFSET we could impinge before losing the lock = +90. This has the effect of increasing our optical gain.

We started using the HPC LOCK IN screen to dither POS on the different BHD SUS. We first started with AS1 (freq = 137.137 Hz, gain = 1000). The sensing matrix element was chosen accordingly (from the demodulated output) and fed to the LO_PHASE; because this affected the AS port alignment this was of course not the best choice. We moved over to LO2 (freq = 318.75 Hz, gain = 1000) but for the longest time couldn't see the dither line at the error point (A-B).

After this we added comb60 notch filters at DCPD_A and DCPD_B input signals. We ended up just feeding the (A-B) error point to LO1, and trying to lock mid fringe, which suceeded without saturation. The gain of the LO_PHASE filter was set to 0.2 (previously 20; attributable to the newly unclipped LO beam intensity?), and again we only enabled FM4 and FM5 for this. After this a dither line at 318.75 Hz finally appeared in the A-B error point! To be continued...

Attachment 1: Screenshot_2022-09-13_17-41-33.png  240 kB  Uploaded Tue Sep 13 19:12:54 2022  | Hide | Hide all
Screenshot_2022-09-13_17-41-33.png
ELOG V3.1.3-